# Quantifying Calibration Error in Modern Neural Networks through Evidence Based Theory

### v1

@@ 4 Problem statement
Traditional metrics like accuracy, precision, and recall focus solely on the performance of neural networks in terms of their output but, fail to capture the uncertainties and subjective aspects associated with their predictions. These metrics do not provide insights into how much confidence one can place in the model’s predictions, particularly in situations where the model might be overconfident or underconfident.
One approach to improving trustworthiness is through confidence calibration, where methods like the Expected Calibration Error (ECE) are used to align predicted probabilities with actual outcomes. However, while ECE provides a useful measure of calibration, it is not easily interpretable and does not offer a comprehensive view of an AI system’s trustworthiness. ECE focuses on the difference between predicted probabilities and actual outcomes but lacks the capability to represent the degrees of belief, disbelief, and uncertainty that are crucial for a more nuanced trust assessment.
Furthermore, existing black-box approaches for quantifying trustworthiness often require access to ground truth data during the operational phase, which can be difficult or even impossible to obtain in real-world scenarios. In some cases, these methods rely on the assumption of an oracle to provide ground truth, but this oracle itself may be unreliable or untrustworthy, further complicating the assessment process. This reliance on unavailable or uncertain ground truth limits the effectiveness of these approaches, highlighting the need for more robust methods that can evaluate trustworthiness without depending on such assumptions.
5 Quantification Framework
This paper tackles the challenges of quantifying the trustworthiness of neural networks by introducing a novel framework grounded in subjective logic. Our primary contribution lies in developing a comprehensive trustworthiness quantification method that extends the traditional Expected Calibration Error (ECE). This framework incorporates subjective measures of belief, disbelief, and uncertainty, offering a more nuanced and interpretable evaluation of model trustworthiness.
The framework operates through the following steps summarized in Algorithm 1:
5.1 Clustering Predicted Probabilities
5.2 Computing Trustworthiness Opinions
For each cluster $i$, we calculate the number of classifications $n_{i}$ that fall into this cluster. We then compute the following:
  • Negative Evidence: This quantity is based on the difference between:
    1. (1)
    2. (2)
      the representative probability ($RP_{i}$).
    The full equation is: $\alpha\mathbbm{1}_{t_{c}>n_{c}\times RP_{i}}(t_{c}-n_{c}\times RP_{i})+\beta\mathbbm{1}_{t_{c}<n_{c}\times RP_{i}}(n_{c}\times RP_{i}-t_{c})$ $\alpha$ and $\beta$ choose based on how we want to penalize over and under-confidence. For simplicity, we use $\alpha=\beta=1$, thus: $\lvert t_{c}-n_{c}\times RP_{i}\rvert=n_{c}\lvert\frac{t_{c}}{n_{c}}-RP_{i}\rvert$.
  • The total number of evidence in our framework is then: $t_{c}+\lvert t_{c}-n_{c}\times RP_{i}\rvert=t_{c}(1+\lvert 1-\frac{n_{c}}{t_{c}}\times RP_{i}\rvert)$, which mostly depends on how far the the predictive probability is from actual probability.
5.3 Fusing Trust Opinions
The trustworthiness opinions for each probability cluster are combined using appropriate fusion operators. This step allows us to synthesize the trust opinions into a single, comprehensive trust opinion that reflects the overall trustworthiness of the neural network’s predictions. The fusion process accounts for the varying levels of trustworthiness across different clusters, ensuring that the final trust opinion is balanced and representative.
@@ 5.3 Fusing Trust Opinions
5.4 From Static to Dynamic
@@
6 Experiments
We evaluate two neural networks, one trained on the MNIST dataset and the other on CIFAR-10. The evaluation consists of assessing the trustworthiness of the models before and after calibration using temperature scaling. To quantify trustworthiness, we applied our algorithm (based on subjective logic) to calculate belief, disbelief, and uncertainty for individual probability clusters, and we used cumulative fusion to combine these opinions. We chose cumulative fusion because of its associativity and commutativity which allows us to generalize this fusion operators to more than two trustworthiness opinion 2.
$&\text{Case I: For }u_{X}^{A}\neq 0\lor u_{X}^{B}\neq 0\\ b_{X}^{(A\diamond B)}(x)&=\frac{b_{X}^{A}(x)u_{X}^{B}+b_{X}^{B}(x)u_{X}^{A}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ u_{X}^{(A\diamond B)}&=\frac{u_{X}^{A}u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ a_{X}^{A\diamond B}(x)&=\frac{a_{X}^{A}(x)u_{X}^{B}+a_{X}^{B}(x)u_{X}^{A}-(a_{X}^{A}(x)+a_{X}^{B}(x))u_{X}^{A}u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}-2u_{X}^{A}u_{X}^{B}},&\text{if }u_{X}^{A}\neq 1\lor u_{X}^{B}\neq 1,\\ a_{X}^{A\diamond B}(x)&=\frac{a_{X}^{A}(x)+a_{X}^{B}(x)}{2},&\text{if }u_{X}^{A}=u_{X}^{B}=1,\\ \\ &\text{Case II: For }u_{X}^{A}=u_{X}^{B}=0\\ b_{X}^{(A\diamond B)}(x)&=\gamma_{X}^{A}b_{X}^{A}(x)+\gamma_{X}^{B}b_{X}^{B}(x),\\ u_{X}^{(A\diamond B)}&=0,\\ a_{X}^{A\diamond B}(x)&=\gamma_{X}^{A}a_{X}^{A}(x)+\gamma_{X}^{B}a_{X}^{B}(x),\\ \\ &\text{where}\\ \gamma_{X}^{A}&=\lim_{u_{X}^{A}\to 0}\frac{u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}},\\ \gamma_{X}^{B}&=\lim_{u_{X}^{B}\to 0}\frac{u_{X}^{A}}{u_{X}^{A}+u_{X}^{B}}.\end{aligned}\right.$ (2)
In our case for binomial opinion, and since we used $W=2$ from Eq. 1 and we do not have an infinite number of evidence we can assume $u_{X}^{A}$ and $u_{X}^{B}\neq 0$. As a result Eq. 2 becomes:
{$b_{X}^{(A\diamond B)}&=\frac{b_{X}^{A}u_{X}^{B}+b_{X}^{B}u_{X}^{A}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ d_{X}^{(A\diamond B)}&=\frac{d_{X}^{A}u_{X}^{B}+d_{X}^{B}u_{X}^{A}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ u_{X}^{(A\diamond B)}&=\frac{u_{X}^{A}u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ a_{X}^{A\diamond B}&=\frac{a_{X}^{A}u_{X}^{B}+a_{X}^{B}u_{X}^{A}-(a_{X}^{A}+a_{X}^{B})u_{X}^{A}u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}-2u_{X}^{A}u_{X}^{B}},&\text{if}u_{X}^{A}\neq 1\lor u_{X}^{B}\neq 1,\\ a_{X}^{A\diamond B}&=\frac{a_{X}^{A}+a_{X}^{B}}{2},&\text{if }u_{X}^{A}=u_{X}^{B}=1,\\ \end{aligned}\right.$ (3)
In our implementation we always keep the number of positive and negative evidence allowing us to just add them and then use Eq.1 to derive the opinion since cumulative fusion exactly does that.
@@ 6 Experiments
6.1 Datasets
We evaluated the models on two well-known datasets:
@@ 6.1 Datasets
  • MNIST: The MNIST dataset consists of 70,000 grayscale images of handwritten digits (0–9), each of size 28×28 pixels. We split the dataset into 60,000 images for training and 10,000 for testing. The task is to classify each image into one of the 10 digit classes.
@@
  • CIFAR-10: The CIFAR-10 dataset contains 60,000 color images of size 32×32, divided into 10 classes, with 6,000 images per class. The dataset is split into 50,000 training images and 10,000 test images. The classes represent various objects, such as airplanes, automobiles, birds, cats, and more. Each image is classified into one of the 10 object categories.
6.2 Neural Network Architectures
@@
We designed two distinct neural network architectures, optimized for each dataset:
  • MNIST Network: A fully connected neural network was employed for MNIST classification. The architecture consists of:
    • The input is a 28×28 grayscale image. This layer flattens the 2D image into a 1D array of 784 elements to serve as the input for the subsequent fully connected layers.
    • A fully connected layer with 128 units, followed by the ReLU activation function. This layer learns 128 features from the input data, allowing the network to model complex relationships in the data.
    • The final fully connected layer with 10 units (one for each digit class). This layer outputs raw scores (logits) for each of the 10 classes. These logits are then converted into probabilities using the softmax function during evaluation and prediction.
    This straightforward architecture is sufficient for classifying MNIST digits due to the simplicity of the dataset.
  • CIFAR-10 Network: A convolutional neural network (CNN) was employed for CIFAR-10 classification. The architecture consists of:
    • Convolutional Layers: The model includes three convolutional layers with ReLU activation. The first layer uses 32 filters, while the subsequent two layers each use 64 filters, all with a 3×3 kernel size. These layers progressively extract low- to high-level features from the input images.
    • Max Pooling Layers: After the first and second convolutional layers, 2×2 max pooling layers are applied to reduce the spatial dimensions of the feature maps, improving computational efficiency.
    • Flatten Layer: The feature maps from the final convolutional layer are flattened into a 1D vector, preparing them for the fully connected layers.
    • Dense Layer: A fully connected layer with 64 units, followed by ReLU activation, enables the model to learn complex patterns from the extracted features.
    • Output Layer: The final layer is a softmax layer with 10 units, providing the probability distribution across the 10 CIFAR-10 classes.
    This architecture is well-suited for the complex task of object classification in CIFAR-10, handling the variations in color, texture, and shapes present in the dataset.
@@
Both neural networks were trained using the Adam optimizer, which adapts learning rates based on the computed gradients, improving convergence speed. The MNIST model was compiled with the Adam optimizer, a sparse categorical crossentropy loss function (with logits). The model was trained for 100 epochs with a batch size of 32, utilizing 90% of the training data for training and 10% as a validation set.
For the CIFAR-10 model, the same Adam optimizer was used, but with sparse categorical crossentropy (without logits). The model was trained for 100 epochs with a batch size of 32, and 10% of the training data was reserved for validation.
6.3 Calibration: Temperature Scaling
@@
After training, we calibrated both models using temperature scaling, a post-processing technique designed to align predicted probabilities with actual outcomes. Temperature scaling modifies the model’s output logits without changing its classification accuracy, thereby improving the reliability of the predicted probabilities.
@@
The calibration process involves an iterative optimization loop where the temperature parameter is fine-tuned over several calibration epochs. During each epoch, the model’s logits are recalibrated by applying the temperature parameter, and the NLL loss is computed between the calibrated probabilities and the true labels. The gradients of the loss are then used to update the temperature parameter, reducing the difference between predicted probabilities and the actual outcomes.
We used $M=10$ groups for computing the trustworthiness. the clusters are $[\frac{i}{10},\frac{i+1}{10}[$ for $i=0,\ldots,8$ with the final cluster covering $[\frac{9}{10},1]$ The representatives are $\frac{i}{10}+\frac{1}{20}$ for $i=0,\ldots,9$.
7 Results and Discussion
@@
This section compares the trustworthiness evaluation results before and after calibration (using temperature scaling) of the MNIST and CIFAR-10 neural networks. The metrics—trust, disbelief, and uncertainty—were assessed over 100 epochs for all labels (0 to 9). The color bar in the figures represents the label corresponding to each curve, allowing us to observe the impact of calibration on individual labels. The results reveal distinct differences between the pre- and post-calibration performance, particularly in how the metrics evolve over time.
@@
7.1 MNIST Results
The results for MNIST model are depicted in Figure 1
  • Trust: Before calibration, trust started at a similar level as post-calibration (around 0.65), but as the training progressed, trust values decreased, fluctuating across different labels. This pattern suggests that the confidence in the network’s probabilities predictions diminished over epochs. In contrast, after calibration, trust remained constant. This demonstrates that temperature scaling helped the network maintain a higher and more consistent level of trust in its probabilities predictions.
  • Disbelief: Before calibration, disbelief started low but increased as the number of training epochs grew, indicating growing overconfidence and/or underconfidence in predictions. However after calibration, disbelief remained constant and stable at around 0.35. This improvement reflects the calibration’s effectiveness in reducing the network’s overconfidence and/or underconfidence in wrong predictions.
  • Uncertainty: Similarly, uncertainty began at a comparable point before and after calibration. However, without calibration, uncertainty tended to increase, showing less evidence during training. Post-calibration, uncertainty remained low and stable, indicating the model’s uncertainty stability.
7.2 CIFAR10 Results
@@
  • Trust: Both pre- and post-calibration trust curves began at similar values (at around 0.7) the increase a little bit for some labels to reach around 0.75 for pre calibration and 0.82 for post calibration. after that without calibration, as training continued, trust decreased to reach 0.48 after 20 epochs and remains constant. After calibration, trust decreases a little to come back to the initial value 0.7 after around 20 epochs and then remained constant. This reflects the impact of calibration in maintaining high confidence in probability predictions.
@@
  • Disbelief: Both pre- and post-calibration disbelief curves began at similar values, around 0.4. Before calibration, disbelief gradually increased as training progressed, reaching a peak of 0.58 after approximately 20 epochs, where it remained relatively constant for the rest of the training. After calibration, disbelief initially increased slightly, but then decreased steadily to return to the initial value of 0.4 around the 20th epoch. It remained stable thereafter. This shows that calibration was effective in controlling the network’s overconfidence in incorrect predictions, leading to more reliable performance over time.
@@
  • Uncertainty: Pre- and post-calibration uncertainty curves also started at similar values, around 0.12. Before calibration, uncertainty increased slightly during the early training epochs, peaking at approximately 0.17 after 20 epochs, and remained at that level for the rest of the training process. After calibration, uncertainty initially increased similarly, but quickly dropped back to the original level of 0.12 around epoch 20 and remained constant thereafter. This demonstrates that calibration helped the model to reduce uncertainty, particularly in later epochs, resulting in more stable and confident predictions.
@@
Overall, the results show that temperature scaling significantly improved the stability of the trustworthiness metrics, maintaining higher trust and lower disbelief and uncertainty, compared to the pre-calibration network, where trust diminished and disbelief and uncertainty increased as training progressed.
7.3 Discussion

### v2

@@ IV Problem statement
Traditional metrics like accuracy, precision, and recall focus solely on the performance of neural networks in terms of their output but fail to capture the uncertainties associated with their predictions. These metrics do not provide insights into how much confidence one can place in the model’s predictions, particularly in situations where the model might be overconfident or underconfident.
One approach to improving trustworthiness is through confidence calibration, where methods like the Expected Calibration Error (ECE) are used to evaluate how well predicted probabilities are aligned with actual outcomes. However, while ECE provides a useful measure of calibration, it is not easily interpretable and does not offer a comprehensive view of an AI system’s trustworthiness. ECE focuses on the difference between predicted probabilities and actual outcomes but lacks the capability to represent the degrees of belief, disbelief, and uncertainty that are crucial for a more nuanced trust assessment.
V Quantification Framework
@@
This paper tackles the challenges of quantifying the trustworthiness of neural networks by introducing a novel framework grounded in subjective logic. Our primary contribution lies in developing a comprehensive trustworthiness quantification method that extends the traditional Expected Calibration Error (ECE). This framework incorporates subjective measures of belief (trust), disbelief (distrust), and uncertainty, offering a more nuanced and interpretable evaluation of model trustworthiness. The framework operates through the following steps summarized in Algorithm 1:
V-A Clustering Predicted Probabilities
V-B Computing Trustworthiness Opinions
For a given class $c$ and for each cluster $i$, we calculate the number of classifications $n_{i}$ that fall into this cluster. We then compute the following:
  • Negative Evidence: This quantity is based on the difference between:
    1. 1.
    2. 2.
      the representative probability ($RP_{i}$).
    The full equation is:
    $\alpha\mathbbm{1}_{t_{c}>n_{c}\times RP_{i}}(t_{c}-n_{c}\times RP_{i})+$
    $\beta\mathbbm{1}_{t_{c}<n_{c}\times RP_{i}}(n_{c}\times RP_{i}-t_{c})$
    $\alpha$ and $\beta$ are chosen based on how much we want to penalize over and under-confidence. For simplicity, we use $\alpha=\beta=1$, thus the negative evidence is:
    $\lvert t_{c}-n_{c}\times RP_{i}\rvert=n_{c}\lvert\frac{t_{c}}{n_{c}}-RP_{i}\rvert$
  • The total number of evidence in our framework is then:
    $t_{c}+\lvert t_{c}-n_{c}\times RP_{i}\rvert=t_{c}(1+\lvert 1-\frac{n_{c}}{t_{c}}\times RP_{i}\rvert)$
    which mostly depends on how far the the predictive probability is from actual probability.
V-C Fusing Trust Opinions
The trustworthiness opinions for each probability cluster are combined using appropriate fusion operators. This step allows us to synthesize the trust opinions into a single, comprehensive trust opinion that reflects the overall trustworthiness of the neural network’s probability predictions for class $c$. The fusion process accounts for the varying levels of trustworthiness across different clusters, ensuring that the final trust opinion is balanced and representative.
The proposed trustworthiness quantification framework is designed to be computationally efficient. Its main steps—clustering predicted probabilities, computing trustworthiness opinions, and fusing them using subjective logic—operate with a complexity of $\mathcal{O}(C\cdot M\cdot N)$ , where $N$ is the number of predictions (i.e. number of data used for the quantification ), $C$ is the number of classes, and $M$ is the number of probability clusters. Since the complexity is linear, the framework scales efficiently and can be integrated into real-world applications without significant computational overhead.
Note that our work shares conceptual similarity with a paper on hallucination detection in LLMs using clustering of output entropy [3]. While that study targets semantic consistency, our method extends ECE with subjective logic to quantify trustworthiness.
V-D From Static to Dynamic
@@
VI Experiments
We evaluate two neural networks, one trained on the MNIST dataset and the other on CIFAR-10. The evaluation consists of assessing the trustworthiness of the models before and after calibration using temperature scaling. To quantify trustworthiness, we apply our algorithm (based on subjective logic) to calculate belief, disbelief, and uncertainty for individual probability clusters, and we use cumulative fusion to combine these opinions. We chose cumulative fusion because of its associativity and commutativity which allow us to generalize this fusion operators to more than two trustworthiness opinion (Equation 2).
{$b_{X}^{(A\diamond B)}&=\frac{b_{X}^{A}u_{X}^{B}+b_{X}^{B}u_{X}^{A}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ d_{X}^{(A\diamond B)}&=\frac{d_{X}^{A}u_{X}^{B}+d_{X}^{B}u_{X}^{A}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ u_{X}^{(A\diamond B)}&=\frac{u_{X}^{A}u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}-u_{X}^{A}u_{X}^{B}},\\ a_{X}^{A\diamond B}&=\frac{a_{X}^{A}u_{X}^{B}+a_{X}^{B}u_{X}^{A}-(a_{X}^{A}+a_{X}^{B})u_{X}^{A}u_{X}^{B}}{u_{X}^{A}+u_{X}^{B}-2u_{X}^{A}u_{X}^{B}},\\ &\text{if }u_{X}^{A}\neq 1\lor u_{X}^{B}\neq 1,\\ a_{X}^{A\diamond B}&=\frac{a_{X}^{A}+a_{X}^{B}}{2},\\ &\text{if }u_{X}^{A}=u_{X}^{B}=1,\\ \end{aligned}\right.$ (2)
In our implementation, for simplicity, we always keep the number of positive and negative evidence allowing us to just add them at the end and then use Eq.1 to derive the opinion since cumulative fusion is just an addition of evidence.
@@ VI Experiments
VI-A Datasets
We evaluate the models on two well-known datasets:
MNIST
@@ VI-A Datasets
The MNIST dataset consists of 70,000 grayscale images of handwritten digits (0–9), each of size 28×28 pixels. We split the dataset into 60,000 images for training and 10,000 for testing. The task is to classify each image into one of the 10 digit classes.
CIFAR-10
@@
The CIFAR-10 dataset contains 60,000 color images of size 32×32, divided into 10 classes, with 6,000 images per class. The dataset is split into 50,000 training images and 10,000 test images. The classes represent various objects, such as airplanes, automobiles, birds, cats, and more. Each image is classified into one of the 10 object categories.
VI-B Neural Network Architectures
@@
We designed two distinct neural network architectures, optimized for each dataset.
For the MNIST model, the input is a 28×28 grayscale image, flattened into a 1D array of 784 elements for the fully connected layers. A fully connected layer with 128 units and ReLU activation extracts 128 features, enabling the network to model complex relationships. The final fully connected layer with 10 units outputs logits for each digit class, which are converted to probabilities using the softmax function. This straightforward architecture effectively classifies MNIST digits due to the dataset’s simplicity.
For CIFAR-10 we use a convolutional neural network (CNN) was used for CIFAR-10 classification. The architecture includes three convolutional layers with ReLU activation: the first uses 32 filters, and the next two use 64 filters, all with 3×3 kernels, to progressively extract features. After the first two convolutional layers, 2×2 max pooling reduces spatial dimensions and improves efficiency. The feature maps from the final convolutional layer are flattened into a 1D vector for the fully connected layers. A dense layer with 64 units and ReLU activation captures complex patterns, and a final softmax layer with 10 units outputs the probability distribution for CIFAR-10 classes. This design effectively handles the dataset’s complexity, including variations in color, and shape.
@@
Both neural networks were trained using the Adam optimizer, improving convergence speed. The MNIST model was compiled with the Adam optimizer, a sparse categorical crossentropy loss function. The model was trained for 100 epochs with a batch size of 32, using 90% of the training data for training and 10% as a validation set.
VI-C Calibration: Temperature Scaling
@@
After training, we calibrate both models using temperature scaling, a post-processing technique designed to align predicted probabilities with actual outcomes. Temperature scaling modifies the model’s output logits without changing its classification accuracy, thereby improving the reliability of the predicted probabilities.
@@
The calibration process involves an iterative optimization loop where the temperature parameter is fine-tuned over several calibration epochs. During each epoch, the model’s logits are recalibrated by applying the temperature parameter, and the NLL loss is computed between the calibrated probabilities and the true labels. The gradients of the loss are then used to update the temperature parameter, reducing the difference between predicted probabilities and actual outcomes.
In the experiments, we used $M=10$ groups for computing the trustworthiness.The clusters are $[\frac{i}{10},\frac{i+1}{10}[$ for $i=0,\ldots,8$ with the final cluster covering $[\frac{9}{10},1]$ The representatives are $\frac{i}{10}+\frac{1}{20}$ for $i=0,\ldots,9$.
VII Results and Discussion
@@
This section compares the trustworthiness evaluation results before and after calibration (using temperature scaling) of the MNIST and CIFAR-10 neural networks. The metrics—trust, distrust, and uncertainty—were assessed over 100 epochs for all labels (0 to 9). The color bar in the figures represents the label corresponding to each curve, allowing us to observe the impact of calibration on individual labels. The results reveal distinct differences between the pre- and post-calibration performance, particularly in how the metrics evolve over time.
@@
VII-A MNIST Results
The results for the MNIST model are depicted in Figure 1.
Trust
Before calibration, trust started at a similar level as post-calibration (around 0.65), but as the training progressed, trust values decreased, fluctuating across different labels. This pattern suggests that the confidence in the network’s probabilities predictions diminished over epochs. In contrast, after calibration, trust remained constant. This demonstrates that temperature scaling helped the network maintain a higher and more consistent level of trust in its probabilities predictions.
distrust
Before calibration, distrust started low but increased as the number of training epochs grew, indicating growing overconfidence and/or underconfidence in predictions. However after calibration, distrust remained constant and stable at around 0.35. This improvement reflects the calibration’s effectiveness in reducing the network’s overconfidence and/or underconfidence in wrong predictions.
Uncertainty
Similarly, uncertainty began at a comparable point before and after calibration. However, without calibration, uncertainty tended to increase, showing less evidence during training. Post-calibration, uncertainty remained low and stable, indicating the model’s uncertainty stability.
VII-B CIFAR10 Results
@@
Both pre- and post-calibration trust curves began at similar values (at around 0.7) then increase a little bit for some labels to reach around 0.75 for pre calibration and 0.82 for post calibration. after that without calibration, as training continued, trust decreased to reach 0.48 after 20 epochs and remains constant. After calibration, trust decreases a little to come back to the initial value 0.7 after around 20 epochs and then remained constant. This reflects the impact of calibration in maintaining high confidence in probability predictions.
distrust
@@
Both pre- and post-calibration distrust curves began at similar values, around 0.4. Before calibration, distrust gradually increased as training progressed, reaching a peak of 0.58 after approximately 20 epochs, where it remained relatively constant for the rest of the training. After calibration, distrust initially increased slightly, but then decreased steadily to return to the initial value of 0.4 around the 20th epoch. It remained stable thereafter. This shows that calibration was effective in controlling the network’s overconfidence in incorrect predictions, leading to more reliable performance over time.
Uncertainty
@@
Pre- and post-calibration uncertainty curves also started at similar values, around 0.12. Before calibration, uncertainty increased slightly during the early training epochs, peaking at approximately 0.17 after 20 epochs, and remained at that level for the rest of the training process. After calibration, uncertainty initially increased similarly, but quickly dropped back to the original level of 0.12 around epoch 20 and remained constant thereafter. This demonstrates that calibration helped the model to reduce uncertainty, particularly in later epochs, resulting in more stable and confident predictions.
@@
Overall, the results show that temperature scaling significantly improved the stability of the trustworthiness metrics, maintaining higher trust and lower distrust and uncertainty, compared to the pre-calibration network, where trust diminished and distrust and uncertainty increased as training progressed.
VII-C Discussion